OpenStudy (anonymous):

Small farm has 70 acres of space to plant.farmer wants to grow corn and green beans. cost $60/acre to grow corn and $30 acre for green beans. Farmer has $1800 to use. Takes farmer 3 days to plant acre of corn and 4days to plant acre of green beans. 120 days available to plant vegetables. Farmer's profit is $180/acre of corn and $100/acre of green beans. How many acres should farmer allow for each vegetable to maximize profit?

OpenStudy (e.cociuba):

Let x = no. of acres for corn Let y = no. of acres for beans : Write an equation for each constraint, put each equation the slope intercept for for graphing : The Acerage equation x + y =< 70 y = -x + 70, (Red) : Cost equation 60x + 30y =< 1800 30y = -60x + 1800 y = -2x + 60, Green : Time equation 3x + 4y =< 120 4y = -3x + 120 y = -.75x + 30 (blue) : Obviously, all values are positive x => 0 y => 0 ; Graph all 3 equations on the same grid : From the graph we can see an intersection at x = 24, y = 12 Check and you can see these values will give us a cost of $1800 and time of 120 days We can ignore the acreage equation, only 36 acres can be planted : "When the farmer sells the vegetables, he can expect to make a profit of $180 per acre for the corn and $100 per acre for the green beans." Profit: 24(180) + 12(100) = 4320 + 1200 = $5520 max profit : But another corner is x=30, y=0; 30 acres of corn and no beans It also will cost $1800 and but only 90 days to plant Profit: 30(180) = $5400, slightly less, but less time required : You can check the 3rd corner. x = 0; y = 30, 30 acres of beans and no corn

OpenStudy (anonymous):

I don't understand how to get constraints for this problem and the steps to solve

OpenStudy (anonymous):

Ok what are the points on the graph and what does the graph look like